a fan is running at 3000 rpm. It is switched off.It comes to rest by uniformly decreasing its angular speed in 10 seconds.The total number of revolution in this period.

Grade:11

3 Answers

Arun
6 years ago
Dear Aadesh
 
thanks= 500 - 250 = 250 rotations.2= 50* 10 + 0.5 (-5) 10\ThetaSo,2t\alphat+ 0.5= w\ThetaNow,2= – 5 rotations / s\alphat,\alpha+oSo, w = 0 = wtime = 10 seconds= 300 rpm = 3000 /60 rotations per second = 50 rpsoGiven , wHi,
 
Regards
Arun (askIITians forum expert)
Arun
6 years ago
thanks= 250 rotations.2= 50* 10 + 0.5 (-5) 10\ThetaSo,2t\alphat+ 0.5= w\ThetaNow,2= – 5 rotations / s\alphat,\alpha+oSo, w = 0 = wtime = 10 seconds= 3000 rpm = 50 rotations per secondoGiven , wHi,
Mohit Jaiswal
5 years ago
3000/6=50 rps . wf=wi+ (alpha)t 
0=50+(alpha)10
-alpla×10=50
Alpha=-5
S=ut+1/2at^2
S=50×10 +1/2×-5×10×10
S=500-250
S=250
So answer is 250 revolutions